OpenStudy (anonymous):

QUESTION 10 OF 10 To borrow money, you pawn your mountain bike. Based on the value of the bike, the pawnbroker loans you $552. One month later, you get the bike back by paying the pawnbroker $851. What annual interest rate did you pay? Round to the nearest tenth of a percent.

OpenStudy (anonymous):

I = Prt, so your interest is $851-$552, P=$552, rate=what you are looking for, and t=1/12 years. solve for r!

OpenStudy (anonymous):

the time is where i am getting confused. I got 299/851 (? =t)

OpenStudy (anonymous):

not sure how to multiply 1/12

OpenStudy (anonymous):

I=Prt, so r=I/(Pt)=(851-552)/(552*(1/12)) you can literally copy and paste (851-552)/(552*(1/12)) into google and it works as a calculator

OpenStudy (anonymous):

multiplying by 1/12 is the same as dividing by 12

OpenStudy (anonymous):

ex: 24*(1/12)=2 24/12=2

OpenStudy (anonymous):

you can check your work with this http://www.webmath.com/simpinterest.html

OpenStudy (anonymous):

That is awesome. It makes it easier when i can see step by step so I can grasp the concept. Thank you so very much!